OpenStudy (anonymous):

what is the LCD of x over 2x+2 = -2x over 4x+4 plus 2x-3 over x+1

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (cruffo):

\[\frac{x}{2x+2} = \frac{-2x}{4x+4} + \frac{2x-3}{x+1}\] Factor the all the denominators first. \[\frac{x}{2(x+1)} = \frac{-2x}{4(x+1)} + \frac{2x-3}{x+1}\] You may want to simplify that second fraction. \[\frac{x}{2(x+1)} = \frac{-x}{2(x+1)} + \frac{2x-3}{x+1}\] Notice that all the fractions have the factor (x+1), and that two of the fractions have 2 as a factor as well. So the LCD = 2(x+1).

OpenStudy (anonymous):

Thank you so much!! Thats so simple and I was making much harder than it was! you are awesome!!!